OpenStudy (icalibear):

Simplify 7x^6y^4 + (x^3y^2)^2

OpenStudy (bibby):

\[\large(x^3y^2)^2=(x^3)^2(y^2)^2 = (x^{3*2}(y^{2*2})\]